# From Heegaard diagrams to surgery

### v2

@@ Abstract.
Mathematics Subject Classification (2020): 57-01, 57K30
Keywords : 3-manifolds, knots, links, Kirby’s calculus
@@ 1. Introduction
By a surface we mean here a closed, connected and orientable surface and a manifold is a closed oriented three-dimensional manifold. Such manifolds are presentable in terms of diagrams in several ways, and we mention here the presentation in terms of Heegaard diagrams and in terms of framed links in $S^{3}$. The latter is usually called surgery presentation. The aim of this note is to provide a precise steps in transforming planar Heegaard diagrams into surgery. The details of a reverse transformation are given, for example, in [1].
We are aware that this procedure is considered as a folklore and that it could be traced back to Lickorish’s results from [7]. However, when we were faced with a concrete problem of finding surgery presentation for some small covers (see [3] or [2] for this notion), a precise algorithm for doing this was not reachable in the literature. A companion to the algorithm provided in this note is a simple example of a small cover obtained by an identification of the facets of eight tetrahedra. This identification gives immediately a (demanding) Heegaard diagram of the obtained manifold (the real projective space), and this diagram could help one to follow the steps of the algorithm. Besides the precision in defining curves and Dehn twists, we ought to point out some heuristics that could considerably shorten the procedure of transforming Heegaard diagrams into surgery and to produce a result of lower complexity.
@@ 2. Heegaard diagrams, surgery and Dehn twists
Let $H$ be a handlebody with $g$ handles. A meridional disk is a properly embedded disk in $H$ (its boundary belongs to the boundary of $H$) such that by removing its neighbourhood from $H$ one obtains a handlebody with $g-1$ handles. A complete system of meridional disks for $H$ consists of $g$ mutually disjoint disks such that by removing their neighbourhoods from $H$ one obtains a ball.
@@
The following remark is analogous to [9, IV, Section 10.1, Theorem 10.2] (see also [8, Remark after Definition 2.3]).
@@
A sufficient condition for $(\Gamma,\alpha,\alpha^{\prime})$, where $\Gamma$ is of genus $g$, to present the sphere $S^{3}$ is that $\Gamma$ is a connected sum $T^{2}_{1}\sharp\ldots\sharp T^{2}_{g}$ of $g$ copies of tori, such that $\alpha_{i}$ and $\alpha^{\prime}_{i}$ are two simple closed curves in $T^{2}_{i}$ that intersect transversally in a single point. This condition is equivalent to: for $i\neq j$, $\alpha_{i}\cap\alpha^{\prime}_{j}=\emptyset$, and for every $i$, $\alpha_{i}$ intersects transversally $\alpha_{i}$ in a single point.
@@
Assume that $\Gamma$, $\alpha$ and $\alpha^{\prime}$ satisfy the conditions of Remark 2.10 and $H$, $H^{\prime}$ are two copies of a handlebody with $g$ handles such that $H$ is standardly embedded in $R^{3}\subset S^{3}$ and the complement of its interior with respect to $S^{3}$ is $H^{\prime}$. The identification of $\Gamma$ with $\partial H=\partial H^{\prime}$ so that $\alpha$ become meridians of $H$, while $\alpha^{\prime}$ become meridians of $H^{\prime}$ is called a proper embedding of $\Gamma$ in $S^{3}$ with respect to $(\alpha,\alpha^{\prime},H,H^{\prime})$.
@@
A surgery is a link in $S^{3}$ whose components are labeled by integers (framings). A surgery is interpreted as a manifold by removing a tubular neighbourhood (a solid torus) of each link component and sewing it back according to the corresponding framing $n$—a meridian of the solid torus is sewed back so that it winds $n$ times in the direction of the meridian and once in the direction of the longitude (see [9]).
@@
Let $l$ be a simple closed curve in an oriented surface $\Gamma$, and let $A$ be an annulus embedded in $\Gamma$ with $l$ as one boundary component. We define intuitively and up to isotopy the right Dehn twist $\theta_{l}\colon\Gamma\to\Gamma$ along $l$. It is the identity in the complement of $A$, and inside $A$ it takes a path transverse to the core of $A$ to a path that wraps once around $A$ turning in the right-hand direction (with respect to the orientation of $\Gamma$). It is evident that $\theta_{l}$ is a homeomorphism and its inverse $\theta^{-1}_{l}$ is called the left Dehn twist along $l$. Both left and right Dehn twists are illustrated in Figure 1.
@@ 3. The algorithm
Let $M$ be presented by a Heegaard diagram $(\Gamma,\alpha,\beta)$ in which a family of curves $\alpha^{\prime}$ is introduced according to Remark 2.10, so that $(\Gamma,\alpha,\alpha^{\prime})$ presents $S^{3}$. The first step in transforming a Heegaard diagram into surgery consists in finding a composition of Dehn twists of $\Gamma$ that maps curves $\alpha^{\prime}_{1},\ldots,\alpha^{\prime}_{g}$ to the curves isotopic to $\beta_{1},\ldots,\beta_{g}$, respectively. This could be done along the lines of the first part of the proof of [7, Theorem 1]. In order to be precise in determining right or left Dehn twists (the right twists have priority in this note) assume that $\Gamma$ is oriented so that its normal is directed towards the handlebody $H^{\prime}$.
@@
  1. (D1)
    If $p$ intersects $q$ in a single point, then $\theta_{p}\circ\theta_{q}$ maps $p$ to a curve isotopic to $q$. (See [7, Lemma 1].)
  2. (D2a)
    If $q$ belongs to a system of attaching circles and $p$ intersects $q$ twice, with zero algebraic intersection, then there is an isotopy moving $p$ away from $q$ and leaving all the other curves from the system fixed. (See the second paragraph of the proof of [7, Lemma 3].)
  3. (D2b)
    Let $A$ and $B$ be two intersection points of $p$ and $q$, which are adjacent on $q$. We denote by $[A,B]$ the segment on $q$ having no other common points with $p$. Assume that $p$ is oriented in the same direction at $A$ and $B$ with respect to $q$ and by walking along $p$ in its direction one has to turn left at $A$ in order to continue the walk along $[A,B]$. Then consider the curve $p_{1}$ that consists of the part of $p$ incoming to $A$, the segment $[A,B]$, and the part of $p$ outgoing from $B$. (See [7, Lemma 2, Case 1].) The curve $\theta_{p_{1}}[p]$ is isotopic to a curve intersecting $q$ in less points than $p$.
  4. (D3)
    Let $A$, $B$ and $C$ be three intersection points of $p$ and $q$, which are adjacent on $q$ and the direction of $p$ induces the circular order $ABC$. We denote by $[A,C]$ the segment on $q$ containing $B$. Assume that $p$ is oriented in alternating directions at $A$, $B$ and $C$ with respect to $q$ and by walking along $p$ in its direction one has to turn left at $A$ in order to continue the walk along $[A,C]$. Then consider the curve $p_{1}$ that consists of the part of $p$ incoming to $A$, the segment $[A,C]$, and the part of $p$ outgoing from $C$. (See [7, Lemma 2, Case 2].) The curve $\theta_{p_{1}}[p]$ is isotopic to a curve intersecting $q$ in less points than $p$.
@@
    We cannot remain forever in (A4) since the procedure from the preceding paragraph must eventually reach the point when $\beta_{t+1}^{\prime}$ intersects no $\beta$-curve. This enables one to reason as in the third paragraph of the proof of [7, Theorem 1] in order to find two curves $l$ and $m$ satisfying the conditions listed in the initial paragraph of (A4).
  5. (A5)
@@
After finding a composition $\theta_{\delta_{n}}\circ\ldots\circ\theta_{\delta_{1}}$ of Dehn twists that maps $\alpha^{\prime}$-curves onto $\beta$-curves, the next step in our procedure is to build a link in $S^{3}$ out of the curves $\delta_{1},\ldots,\delta_{n}$. This link consists of parallel copies of $\delta_{1},\ldots,\delta_{n}$ in the handlebody $H$. This is done by making a parallel copy of $\Gamma$ shallow inside $H$ (or “digging the trench”), layer by layer, so that $\delta_{n}$ is the deepest. In order to calculate the framing of each component of the link we rely on the following remark and lemma.
@@
Let us consider the (right) Dehn twist $\theta_{l}$ with respect to a curve $l$ on $\Gamma$ whose parallel copy in $H$ makes a component $\mu$ of the link. Let $k\in\mathbb{Z}$ be the self-linking of $l\subseteq\Gamma$, i.e. the linking number of $l$ and $\mu$, assuming they are codirected. Then the framing of $\mu$ is $k+1$. (It should be noted that this $+1$ comes from the fact that $\theta_{l}$ is a right Dehn twist with respect to the orientation of $\Gamma$ whose normal is directed outside of $H$—if one switches from right to left twists, i.e. to $\theta_{l}^{-1}$, then the framing of $\mu$ is $k-1$, but there is no need for this here.)
For the following lemma we assume that for every $1\leq j\leq g$, the curves $\alpha_{j}$ and $\alpha^{\prime}_{j}$ are oriented so that the intersection number $i_{A}(\alpha_{j},\alpha^{\prime}_{j})$ is $-1$. Moreover, the embedding of $\Gamma$ is proper with respect to $(\alpha,\alpha^{\prime},H,H^{\prime})$.
@@
Note that the self-linking of $l$ is equal to the linking number of $l$ and its parallel, codirected copy in $\Gamma$. Let us assume that our planar Heegaard diagram looks like as in the picture below (which is possible according to Remark 2.10 with a help of isotopy). We can also assume that $l$ meets $\alpha$ and $\alpha^{\prime}$-curves transversely.
Note that apart of neighbourhoods of the pairs of curves $\alpha_{j}$, $\alpha^{\prime}_{j}$ there is no contribution to the self-linking number of $l$. We can observe each of these neighbourhoods separately. Take a look at the leftmost picture in Figure 2, which illustrates a neighbourhood of $\alpha_{j}$ and $\alpha^{\prime}_{j}$ and a pair of intersection points—one of $l$ and $\alpha_{j}$ and the other of $l$ and $\alpha^{\prime}_{j}$ (negative and positive intersection points are marked in black and red, respectively). It is evident that this pair of intersection points contributes by $-1$ to the sum (3.1).
Refer to caption Refer to caption Refer to caption
Figure 2.
On the other hand, this pair contributes to the self-linking of $l$ again by $-1$, which is justified by the picture in the middle of Figure 2 (parallel, codirected copies are added to the diagram as well as a “bridge” connecting two copies of the intersection point of $l$ and $\alpha_{j}$, which is “above” the diagram). The rightmost picture in Figure 2 shows that the orientation of $l$ does not affect the score. By summing over all these pairs, we get the desired result. ∎
@@
  2. (S2)
    Erase the bottom row of $\alpha$-curves and all the curves from the diagram not involved in $\delta$. Then introduce “bridges” for every curve intersecting $\alpha$-curves.
  3. (S3)
@@
  4. (S4)
    Make parallel copies of the curves involved in $\delta$. This is done step by step starting from $\delta_{2}$.

### v3

@@ Abstract.
Corresponding author.
00footnotetext: Mathematics Subject Classification (2020): 57-01, 57K30.00footnotetext: Key words and phrases: 3-manifolds, Kirby’s calculus, Heegaard diagram, surgery presentation.
@@ 1. Introduction
By a surface we mean here a closed, connected and orientable surface and a manifold is a closed oriented three-dimensional manifold. (Homeomorphisms between manifolds are orientation preserving.) Such manifolds are presentable in terms of diagrams in several ways, and we mention here the presentation in terms of Heegaard diagrams and in terms of framed links in $S^{3}$. The latter is usually called surgery presentation. The aim of this note is to provide a precise steps in transforming planar Heegaard diagrams into a surgery. The details of a reverse transformation are given, for example, in [1].
We are aware that this procedure is considered as a folklore and that it could be traced back to Lickorish’s results from [8]. However, when we were faced with a concrete problem of finding surgery presentation for some small covers (see [3] or [2] for this notion), a precise algorithm for doing this was not reachable in the literature. A companion to the algorithm provided in this note is a simple example of a small cover obtained by an identification of the facets of eight tetrahedra. This identification gives immediately a (demanding) Heegaard diagram of the obtained manifold (the real projective space), and this diagram could help one to follow the steps of the algorithm. Besides the precision in defining curves and Dehn twists, we ought to point out some heuristics that could considerably shorten the procedure of transforming Heegaard diagrams into a surgery and to produce a result of lower complexity.
@@ 2. Heegaard diagrams, surgery and Dehn twists
Let $H$ be a handlebody with $g$ handles. A meridional disk is a properly embedded disk in $H$ (its boundary belongs to the boundary of $H$) such that by removing its neighbourhood from $H$ one obtains a handlebody with g1g-1 handles. A complete system of meridional disks for $H$ consists of $g$ mutually disjoint disks such that by removing their neighbourhoods from $H$ one obtains a ball.
@@
The following remark is analogous to [11, IV, Section 10.1, Theorem 10.2] (see also [10, Remark after Definition 2.3]).
@@
A sufficient condition for $(\Gamma,\alpha,\alpha^{\prime})$, where $\Gamma$ is of genus $g$, to present the sphere $S^{3}$ is that $\Gamma$ is a connected sum $T^{2}_{1}\sharp\ldots\sharp T^{2}_{g}$ of $g$ copies of tori, such that $\alpha_{i}$ and $\alpha^{\prime}_{i}$ are two simple closed curves in $T^{2}_{i}$ that intersect transversally in a single point. This condition is equivalent to: for $i\neq j$, $\alpha_{i}\cap\alpha^{\prime}_{j}=\emptyset$, and for every $i$, $\alpha_{i}$ intersects transversally $\alpha^{\prime}_{i}$ in a single point.
@@
Assume that $\Gamma$, $\alpha$ and $\alpha^{\prime}$ satisfy the conditions of Remark 2.10 and $H$, $H^{\prime}$ are two copies of a handlebody with $g$ handles such that $H$ is standardly embedded in $\mathbb{R}^{3}\subset S^{3}$ and the complement of its interior with respect to $S^{3}$ is $H^{\prime}$. The identification of $\Gamma$ with $\partial H=\partial H^{\prime}$ so that $\alpha$ become meridians of $H$, while $\alpha^{\prime}$ become meridians of $H^{\prime}$ is called a proper embedding of $\Gamma$ in $S^{3}$ with respect to $(\alpha,\alpha^{\prime},H,H^{\prime})$.
@@
A surgery is a link in $S^{3}$ whose components are labeled by integers (framings). A surgery is interpreted as a manifold by removing a tubular neighbourhood (a solid torus) of each link component and sewing it back according to the corresponding framing nn—a meridian of the solid torus is sewed back so that it winds $n$ times in the direction of the meridian and once in the direction of the longitude (see [11]).
@@
Let $l$ be a simple closed curve in an oriented surface $\Gamma$, and let $A$ be an annulus embedded in $\Gamma$ with $l$ as one boundary component. We define intuitively and up to isotopy the right Dehn twist $\theta_{l}\colon\Gamma\to\Gamma$ along $l$. It is the identity in the complement of $A$, and inside $A$ it takes a path transverse to the core of $A$ to a path that wraps once around $A$ turning in the right-hand direction (with respect to the orientation of $\Gamma$). It is evident that $\theta_{l}$ is a homeomorphism and its inverse $\theta^{-1}_{l}$ is called the left Dehn twist along $l$. Both left and right Dehn twists are illustrated in Figure 1 (see [9, Introduction] for a more formal definition of [Dehn] twist).
@@ 3. The algorithm
Let $M$ be presented by a Heegaard diagram $(\Gamma,\alpha,\beta)$ in which a family of curves $\alpha^{\prime}$ is introduced according to Remark 2.10, so that $(\Gamma,\alpha,\alpha^{\prime})$ presents $S^{3}$. Consider a proper embedding of $\Gamma$ with respect to $(\alpha,\alpha^{\prime},H,H^{\prime})$.
The first step in transforming this Heegaard diagram into a surgery for $M$ consists in finding a composition of Dehn twists of $\Gamma$ that maps curves $\alpha^{\prime}_{1},\ldots,\alpha^{\prime}_{g}$ to the curves isotopic to $\beta_{1},\ldots,\beta_{g}$, respectively. This could be done along the lines of the first part of the proof of [8, Theorem 1]. In order to be precise in determining right or left Dehn twists (the right twists have priority in this note) assume that $\Gamma$ is oriented so that its normal is directed towards the handlebody $H^{\prime}$.
@@
  • (D1)
    If $p$ intersects $q$ in a single point, then $\theta_{p}\circ\theta_{q}$ maps $p$ to a curve isotopic to $q$. (See [8, Lemma 1].)
  • (D2a)
    If $q$ belongs to a system of attaching circles and $p$ intersects $q$ twice, with zero algebraic intersection, then there is an isotopy moving $p$ away from $q$ and leaving all the other curves from the system fixed. (See the second paragraph of the proof of [8, Lemma 3].)
  • (D2b)
    Let $A$ and $B$ be two intersection points of $p$ and $q$, which are adjacent on $q$. We denote by $[A,B]$ the segment on $q$ having no other common points with $p$. Assume that $p$ is oriented in the same direction at $A$ and $B$ with respect to $q$ and by walking along $p$ in its direction one has to turn left at $A$ in order to continue the walk along $[A,B]$. Then consider the curve $p_{1}$ that consists of the part of $p$ incoming to $A$, the segment $[A,B]$, and the part of $p$ outgoing from $B$. (See [8, Lemma 2, Case 1].) The curve $\theta_{p_{1}}[p]$ is isotopic to a curve intersecting $q$ in less points than $p$.
  • (D3)
    Let $A$, $B$ and $C$ be three intersection points of $p$ and $q$, which are adjacent on $q$ and the direction of $p$ induces the circular order $ABC$. We denote by $[A,C]$ the segment on $q$ containing $B$. Assume that $p$ is oriented in alternating directions at $A$, $B$ and $C$ with respect to $q$ and by walking along $p$ in its direction one has to turn left at $A$ in order to continue the walk along $[A,C]$. Then consider the curve $p_{1}$ that consists of the part of $p$ incoming to $A$, the segment $[A,C]$, and the part of $p$ outgoing from $C$. (See [8, Lemma 2, Case 2].) The curve $\theta_{p_{1}}[p]$ is isotopic to a curve intersecting $q$ in less points than $p$.
@@
    We cannot remain forever in (A4) since the procedure from the preceding paragraph must eventually reach the point when $\beta_{t+1}^{\prime}$ intersects no $\beta$-curve. This enables one to reason as in the third paragraph of the proof of [8, Theorem 1] in order to find two curves $l$ and $m$ satisfying the conditions listed in the initial paragraph of (A4).
  • (A5)
@@
After finding a composition $\theta_{\delta_{n}}\circ\ldots\circ\theta_{\delta_{1}}$ of Dehn twists that maps $\alpha^{\prime}$-curves onto $\beta$-curves, the next step in our procedure is to build a link in $S^{3}$ out of the curves $\delta_{1},\ldots,\delta_{n}$. This link consists of parallel copies of $\delta_{1},\ldots,\delta_{n}$ in the handlebody $H$ (see the proof of Proposition 3.3). This is done by making a parallel copy of $\Gamma$ shallow inside $H$ (or “digging the trench”), layer by layer, so that $\delta_{n}$ is the deepest. In order to calculate the framing of each component of the link we rely on the following remark (that stems from the proof of Proposition 3.3) and lemma.
@@
Let us consider the (right) Dehn twist $\theta_{l}$ with respect to a curve $l$ on $\Gamma$ whose parallel copy in $H$ makes a component $l^{\prime}$ of the link. Let $k\in\mathbb{Z}$ be the self-linking of $l\subseteq\Gamma$, i.e. the linking number of $l$ and $l^{\prime}$, assuming they are codirected. Then the framing of $l^{\prime}$ is $k+1$. (It should be noted that this +1+1 comes from the fact that $\theta_{l}$ is a right Dehn twist with respect to the orientation of $\Gamma$ whose normal is directed outside of HH—if one switches from right to left twists, i.e. to $\theta_{l}^{-1}$, then the framing of $l^{\prime}$ is k1k-1, but there is no need for this here.)
For the following lemma we assume that for every $1\leq j\leq g$, the curves $\alpha_{j}$ and $\alpha^{\prime}_{j}$ are oriented so that the intersection number $i_{A}(\alpha_{j},\alpha^{\prime}_{j})$ is 1-1. Moreover, the embedding of $\Gamma$ is proper with respect to $(\alpha,\alpha^{\prime},H,H^{\prime})$.
@@
Note that the self-linking of $l$ is equal to the linking number of $l$ and its parallel, codirected copy in $\Gamma$. Let us assume that our planar Heegaard diagram looks like one in Figure 2 (which is possible according to Remark 2.10 with a help of an isotopy). We can also assume that $l$ meets $\alpha$ and $\alpha^{\prime}$-curves transversely.
Figure 2. Standard planar Heegaard diagram of $S^{3}$
Note that apart of neighbourhoods of the pairs of curves $\alpha_{j}$, $\alpha^{\prime}_{j}$ there is no contribution to the self-linking number of $l$. We can observe each of these neighbourhoods separately. Take a look at the leftmost picture in Figure 3, which illustrates a neighbourhood of $\alpha_{j}$ and $\alpha^{\prime}_{j}$ and a pair of intersection points—one of $l$ and $\alpha_{j}$ and the other of $l$ and $\alpha^{\prime}_{j}$ (negative and positive intersection points are marked in black and red, respectively). It is evident that this pair of intersection points contributes by 1-1 to the sum (3.1).
Figure 3. Curve $l$ meets the $j$-th handle
On the other hand, this pair contributes to the self-linking of $l$ again by 1-1, which is justified by the picture in the middle of Figure 3 (parallel, codirected copies are added to the diagram as well as a “bridge” connecting two copies of the intersection point of $l$ and $\alpha_{j}$, which is “above” the diagram). The rightmost picture in Figure 3 shows that the orientation of $l$ does not affect the score. By summing over all these pairs, we get the desired result. ∎
@@
  • (S2)
    Erase one copy of each $\alpha$-curve (assuming that $\alpha$ and $\alpha^{\prime}$-curves are positioned as in Figure 2, erase the bottom row of the copies of $\alpha$-curves). Moreover, erase all the curves from the diagram not involved in the composition $\delta$. Then introduce “bridges” for every curve intersecting $\alpha$-curves. Bridges are disjoint arcs connecting the two endpoints of curves, which become arcs once we delete one of the $\alpha$-curves. These arcs are “above” every other curve in the configuration.
  • (S3)
@@
  • (S4)
    The family of curves obtained by the previous steps does not represent a projection of a link since some intersections are not transversal and except for the “bridges”, under and over crossings are not determined. Our (S4) transforms, step by step, this family of curves into a framed link: fix the curve $\delta_{1}$ and make a parallel copy of $\delta_{2}$ in $H$. This completely determines the under and over crossings of $\delta_{2}$ with the other curves present at the picture (see the left hand side of Figure 8, where $\delta_{1}$ and $\delta_{2}$ are denoted by 1 and 2, respectively). Then repeat this with $\delta_{3}$ and so on. Note that in some cases the curves obtained by (S1)-(S3) produce multiple parallel copies. The example in Section 4 provides a detailed illustrations of this step.
The following proposition is just a reformulation of [8, Theorem 2].
Proposition 3.3.
The surgery obtained by our algorithm results in a manifold homeomorphic to the manifold presented by the Heegaard diagram.
Proof.
Let $M$ be the manifold presented by a Heegaard diagram $(\Gamma,\alpha,\beta)$. Consider a proper embedding of $\Gamma$ with respect to $(\alpha,\alpha^{\prime},H,H^{\prime})$ and let $h\colon\Gamma\to\Gamma$ be an orientation preserving homeomorphism that for every $1\leq i\leq g$ maps $\alpha^{\prime}_{i}$ to a curve isotopic to $\beta_{i}$. Our manifold $M$ is
$H\cup_{h}H^{\prime}=((H\times\{0\})\cup(H^{\prime}\times\{1\}))/\sim,$
where $\sim$ is such that for every $x\in\Gamma$ we have $(h(x),0)\sim(x,1)$.
Figure 4. A cross section of the handlebody $H$ normal to the curve $l$
Assume for a moment that $h$ is just a Dehn twist $\theta_{l}$. Let $l^{\prime}$ be a parallel copy of $l$ shallow inside $H$. Make a tubular neighbourhood $\mathbb{T}$ of $l^{\prime}$ and remove it from $H$ (see Figure 4). One can easily extend $h$ to a homeomorphism $\chi\colon H-\mathbb{T}\to H-\mathbb{T}$, which is the identity in $H-\mathbb{T}$, except in the shaded region above $\mathbb{T}$ (see Figure 4), where it is defined in every layer as $\theta_{l}$. By relying on a straightforward topological argument (see for example [4, Corollary 2.2]) one concludes that $(H-\mathbb{T})\cup_{h}H^{\prime}$ is homeomorphic to
$(H-\mathbb{T})\cup_{\chi^{-1}\circ h}H^{\prime}=(H-\mathbb{T})\cup H^{\prime}=S^{3}-\mathbb{T}.$
All that remains is to glue back the solid torus $\bar{\mathbb{T}}$ at the right position. For this, only important thing is how we glue a neighbourhood of a meridional disk of $\bar{\mathbb{T}}$. Assume that the blue square in Figure 4 is the meridian of $\bar{\mathbb{T}}$. Then its left, right, and bottom side could be glued back by identity, while the top side must be glued to the path resembling the image under the left Dehn twist of a radius of the annulus in Figure 1 (see Figure 5). (This switch from right to left Dehn twist may confuse the reader—precise and formal arguments are given in [4, Section 2].)
Figure 5. Gluing the solid torus $\bar{\mathbb{T}}$ back following the attaching curve $\mu$
The left-hand side of Figure 6 illustrates the curve $l^{\prime}$ (black), the attaching curve $\mu$ (blue) for the meridian and the red curve (the trace of the red dot in Figure 4), which is parallel both to $l$, and $l^{\prime}$. At the right-hand side of Figure 6 the case when $\theta_{l}$ is replaced by $\theta_{l}^{-1}$ is illustrated. In the case of $\theta_{l}$, we conclude that the linking number of $l^{\prime}$ and $\mu$ is by 1 greater than the self-linking of $l$. This linking number is the framing of the component $l^{\prime}$ of the surgery.
Figure 6. Winding of the attaching curve
In the case when $h$ is a composition of several Dehn twists, we just repeat the above procedure save that we start from the rightmost Dehn twist, and the parallels of other curves are chosen deeper and deeper in $H$ (see [4, paragraphs preceding Proposition 2.7] for a detailed explanation.) ∎